OpenStudy (anonymous):

solve equation by factoring 8x^2+21=-59x what could you factor besides the x's

OpenStudy (anonymous):

f(x)=8x^2 +59x +21=0 x's=[-59+&-(59^2-32x21)}/16=-3/8 & -7 f(x)=8(x+3/8)(x+7)=(8x+3)(x+7)

OpenStudy (anonymous):

write the eqn as 8x^2 + 59x + 21=0

OpenStudy (cwrw238):

8x^2 + 59x + 21 = 0 factoring 8 = 2*2*2 21 = 3 * 7 7 * 8 = 56 3 * 1 = 3 adding = 59 - factors are what abayomi12 said

OpenStudy (anonymous):

ookkkk... i think i get it